Graph each system of equations using the coordinate plane provided. Then determine whether the system has one solution, no solution or infinitely many solutions. If the system has one solution, name it.
x+3y=3
y=-1/3x+3

Answer by stanbon(57384)

x+3y=3
y=-(1/3)x+3
--------------
1st: x+3y=3
2nd: x+3y=9
----------------
The two lines are parallel.
There is no solution.
The system is inconsistent.
Cheers,
Stan H.